A mechanic tightens the lugs on a tire by applying a torque of 100 J at an angle of 80° to the line of action.

What force is applied if the wrench is 0.3 meters long? How long must the wrench be if the mechanic is only

capable of applying a force of 180N?


1
Expert's answer
2021-03-26T11:38:50-0400

"\tau=Fl\sin\alpha\to F=\frac{\tau}{l\sin\alpha}=\frac{100}{0.3\cdot\sin80°}=338.5\ (N)"


"l=\frac{\tau}{F\sin\alpha}=\frac{100}{180\cdot\sin80°}=0.564\ (m)" . Answer